OpenStudy (anonymous):

the senate has 17 women senators. If you randomly select 3 different senators to visit, what is the probability that: a) all of the senators would be women? b) two of the senators would be women? c)at least one of the senators would be a woman?

OpenStudy (anonymous):

looked good to me

OpenStudy (anonymous):

you can use \[\frac{17}{100}\times \frac{16}{99}\times \frac{15}{98}\] or \[\frac{\binom{17}{3}}{\binom{100}{3}}\] you will get the same answer

OpenStudy (anonymous):

first method is quicker second method generalizes better for example 2, female, one male \[\frac{\binom{17}{2}\binom{83}{1}}{\binom{100}{3}}\]